OpenStudy (anonymous):

Solve the system by the substitution method. Show your work. 2y - x = 5 x2 + y2 - 25 = 0

OpenStudy (anonymous):

solve for x in the first equation first. 2y - x = 5 add x to both sides and subtract 5 from both sides 2y - 5 = x x= 2y - 5 now substitute this value into the second equation and solve for y.

OpenStudy (anonymous):

(2y - 5)^2 + y^2 - 25 = 0 4y^2 - 20y + 25 + y^2 - 25 = 0 5y^2 - 20y = 0 5y(y - 4)=0 now solve this for y.

OpenStudy (anonymous):

u will get y=0 and y=4 now sustitute these y-values into the equation x=2y - 5 and solve for x. u should get two pairs of solutions

OpenStudy (anonymous):

thanks i got it already!

OpenStudy (anonymous):

np :)